OpenStudy (anonymous):

what the answer (8+j4)(6+9j)

OpenStudy (anonymous):

it is a complex number alisha.

jimthompson5910 (jim_thompson5910):

since j is imaginary, j^2 = -1, so 36j^2 = -36. I'll let you take it from there.

OpenStudy (anonymous):

yes it is. can you help me to prove absolte of complex number? for example \[\left| 2x+5j \right|>2\]

jimthompson5910 (jim_thompson5910):

well |a+bi| = sqrt(a^2+b^2), so that might help.